In an intelligence test administered on 1000 children, the average was 60 and the standard deviation was 20. Assuming that the marks obtained by the children follow a normal distributed, find the number of children who have scored (i) over 90 marks, (ii) below 40 marks and (iii) between 50 and 80 marks. Given that P(0 < Z < 1.5) = 0.4332; P(Z < 1) = 0.8413; P(Z < 0.5) = 0.6915. 


1
Expert's answer
2022-03-02T08:08:45-0500

(i)


"P(X>90)=1-P(Z\le\dfrac{90-60}{20})"

"=1-P(Z\le1.5)=1-(0.5+0.4332)=0.0668"



"0.0668(1000)=67"


(ii)


"P(X<40)=P(Z<\dfrac{40-60}{20})=P(Z<-1)"

"=1-P(Z<1)=1-0.8413=0.1587"



"0.1587(1000)=159"



(iii)


"P(50<X<80)=P(Z<\dfrac{80-60}{20})"

"-P(Z\le\dfrac{50-60}{20})=P(Z<1)-P(Z\le-0.5)"

"=0.8413-(1-0.6915)=0.5328"



"0.5328(1000)=533"